OpenStudy (anonymous):

If one number is 33 1/3 % bigger than the other, then how many % is the other number smaller than the first one?

OpenStudy (anonymous):

a = (1 + 1/3)b b = a / (1 + 1/3) b = (3/4)a b = a - (1/4)a so 25% less

OpenStudy (anonymous):

Thanks!